OpenStudy (anonymous):

Find the range: f(x)=2^x+1 g(x)=-2^x+1

OpenStudy (lukecrayonz):

\[limf(x)\rightarrow 0 as x \rightarrow \infty\]

Directrix (directrix):

The range of f(x) = 2 ^ x + 1 is the set of positve real numbers greater than one. In interval notation, (1, ∞). The range of g(x) = -2^x + 1 is the set of negative real number less than one. In interval notation, (1, -∞). http://jwilson.coe.uga.edu/emt668/EMAT6680.2002/Jackson/EMAT%206000/topic%208/topic%208-exp%20func.html